### v1

# Adaptive Federated Learning with Functional Encryption: A Comparison of Classical and Quantum-safe Options

Abstract
Federated Learning (FL) is a collaborative method for training machine learning models while preserving the confidentiality of the participants’ training data. Nevertheless, FL is vulnerable to reconstruction attacks that exploit shared parameters to reveal private training data.
In this paper, we address this issue in the cybersecurity domain by applying Multi-Input Functional Encryption (MIFE) to a recent FL implementation for training ML-based network intrusion detection systems. We assess both classical and post-quantum solutions in terms of memory cost and computational overhead in the FL process, highlighting their impact on convergence time.
Index Terms:
Federated Learning, Functional Encryption, Cybersecurity, Network Intrusion Detection, DDoS attacks
I Introduction
Federated Learning (FL) is a recent paradigm for training Machine Learning (ML) models with distributed datasets. The key property of FL is that it allows a federation of multiple parties to train a common model without requiring them to share the training data. FL was initially proposed and tested for image classification tasks [1]. However, it has recently been adopted in other domains, such as cybersecurity [2, 3].
In the cybersecurity context, FLAD (adaptive Federated Learning Approach to DDoS attack detection) [2, 4] addresses some of the limitations of FL specific to this domain. Specifically, adaptive Federated Learning Approach to DDoS attack detection (FLAD) introduces a mechanism in which participants (also called “clients”) share the accuracy score of the global model, as measured on their local validation sets, with the FL orchestrator (also called “server”). This approach enables FLAD to outperform FedAVG, the algorithm at the core of FL, in both accuracy and training efficiency on unbalanced and heterogeneous datasets of network intrusions. In addition, the server can use these accuracy scores to implement a stopping criterion to terminate the FL once predefined target conditions are met (e.g., an average accuracy threshold).
Although both FedAVG and FLAD aim at preserving the confidentiality of clients’ training data, clients in both cases are exposed to reconstruction attacks carried out by a malicious server. Such attacks can infer details of the client’s original training data [5] by exploiting the knowledge of the global model’s architecture and information shared by the clients. This includes the global model’s parameters (weights and biases) and gradients, number of training samples in the case of FedAVG, or local validation accuracy in the case of FLAD.
To tackle this issue, we introduce Functional Encryption (FE) techniques into FL for cybersecurity applications by integrating Multi-Input (Inner Product) Functional Encryption (MIFE) with FLAD. MIFE allows the FL server to execute partial computations on encrypted data across multiple inputs, without revealing the underlying plaintext. With FLAD, the multiple inputs are represented by the clients’ parameters and the accuracy scores computed by the clients on their local datasets and used to configure personalised training epochs and steps. As the server only operates on encrypted data, the clients’ original training data remain protected from reconstruction attacks.
The remainder of this paper is structured as follows: Section II introduces the MIFE schemes evaluated in this work. Section III reviews and discusses the related work. Section IV provides a threat model analysis. Section V outlines the methodology used to apply the MIFE schemes to FLAD in order to prevent reconstruction attacks. Section VI and VII detail the experimental setup and results respectively. Finally, the conclusions are given in Section VIII.
II Background
Functional Encryption [6] is an extended form of the public-key setting, where a functional key $sk_{f}$ can be derived from a master secret key $msk$ for a certain function $f$. By applying $sk_{f}$ to the encryption of a plaintext $x$, $f(x)$ will be revealed without decrypting the ciphertext.
A single-input Inner Product Functional Encryption (IPFE) is a type of FE that supports the evaluation of inner product on the encrypted data. Given the encryption of a plaintext vector $\mathbf{x}=({x}_{1},\dots,{x}_{m})$, and a functional key $sk_{y}$ linked with a vector $\mathbf{y}=({y}_{1},\dots,{y}_{m})$, the decryption function, run with $sk_{y}$, outputs $\langle\mathbf{x},\mathbf{y}\rangle=\sum_{i=1}^{m}{x}_{i}{y}_{i}$.
The single-input IPFE can be lifted in a multi-input (MI) setting where different vectorial plaintexts $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\in\mathbb{Z}_{q}^{m}$ are encrypted with different keys and in the decryption phase, using a functional key $sk_{y}$ linked to a vector $\mathbf{y}=(\mathbf{y}_{1},\dots,\mathbf{y}_{n})$, the inner product (Eq. 1) will be revealed.
$\sum_{i=1}^{n}{\langle{\bf x}_{i},{\bf y}_{i}\rangle}$ (1)
The IPFE schemes are composed by four algorithms. A Setup phase in which the public parameters, the master secret key and the clients’ private keys are generated. A KeyGen phase in which the functional key, linked with a vector $\mathbf{y}$, is derived from the master secret key. An Encryption phase where the plaintexts are encrypted and a Decryption phase where the inner product is computed using the functional key.
After a comprehensive review of the literature we have selected two classes of MIFE schemes for inner product: the first one bases its security on the Decisional Diffie Hellman (DDH) problem, while the second is constructed over the Learning With Errors (LWE) problem. While the first one is already used in some construction for the Federated Learning [7], we have not found any benchmark in the second setting, at the best of our knowledge, even if it is a post-quantum solution.
Algorithm 1 A brief description of the DDH-based single-input IPFE algorithm in the selective-secure (left) and adaptive-secure (right) setting.
procedure Setup($1^{\lambda},m$)
     $(\mathcal{G},q,\text{g})\leftarrow\textsc{GroupGen}(1^{\lambda})$
$\mathbf{s}\leftarrow_{R}\mathbb{Z}_{q}^{m}$ $a\leftarrow_{R}\mathbb{Z}_{q},\ \mathbf{a}=(1,a)^{\top}$
$\mathbf{h}\leftarrow(\text{g}^{s_{1}},\dots,{\text{g}}^{s_{m}})$ $\textbf{W}\longleftarrow_{R}\mathbb{Z}_{q}^{m\times 2}$
[…]
procedure Enc($mpk,\mathbf{x}\in\mathbb{Z}_{q}^{m}$)
     $r\leftarrow_{R}\mathbb{Z}_{q}$
$ct_{0}\leftarrow\text{g}^{r}$ $\mathbf{ct}^{\prime}\leftarrow\text{g}^{\mathbf{a}r}$
$\forall i\in[m]\ ct_{i}\leftarrow h_{i}^{r}\ \text{g}^{x_{i}}$ $\mathbf{ct}^{\prime\prime}\leftarrow\text{g}^{\mathbf{x}+\mathbf{W}\mathbf{a}{r}}$
[…]
     return $\mathbf{ct}$
procedure KeyGen($msk,\mathbf{y}\in\mathbb{Z}_{q}^{m}$)
$sk_{\mathbf{y}}\leftarrow(\langle\mathbf{y},\mathbf{s}\rangle,\ \mathbf{y})$ $sk_{\mathbf{y}}\leftarrow(\ \mathbf{W}^{\top}\mathbf{y},\mathbf{y})$
     return $sk_{\mathbf{y}}$
procedure Dec($\mathbf{ct},\ sk_{\mathbf{y}}=(d,\mathbf{y})$)
$C\leftarrow\dfrac{\prod_{i=1}^{m}ct_{i}^{y_{i}}}{ct_{0}^{d}}$ $C\leftarrow\frac{\prod_{i=1}^{m}({ct^{\prime\prime}}_{i})^{y_{i}}}{\prod_{i=1}^{2}({ct^{\prime}}_{i})^{d_{i}}}$
     $res\leftarrow\textsc{Dlog}_{\text{g}}(C)$
     return $res$
In both cases, in order to design a multi-input scheme we start from a single-input scheme and lift it to a multi-input setting through the compiler presented in [8]. To cover a wider range of use cases, we use both a selective-secure and an adaptive-secure [9] single-input scheme for each problem, thus obtaining different schemes with varying properties. In the following, we are going to use the colors above to stress the diffences in the selective and adaptive case.
II-1 DDH-based FE
both the selective-secure and the adaptive-secure scheme are based on the plain DDH problem. The first one is derived from the ElGamal PKE and translated into an IPFE scheme [10], while the second is discussed in detail in [8]. Let GroupGen be a probabilistic polynomial time algorithm that takes as input a security parameter $1^{\lambda}$ and outputs $(\mathcal{G},\text{q},\text{g})$, where $\mathcal{G}$ is a group of prime order q generated by g. Algorithm 1 briefly describes the schemes.
Algorithm 2 A brief description of the LWE-based single-input IPFE algorithm in the selective-secure (left) and adaptive-secure (right) setting.
procedure Setup($1^{\lambda},m$)
     $\mathbf{A}\longleftarrow_{R}\mathbb{Z}_{q}^{M\times N}$
$\mathbf{S}\longleftarrow_{R}\mathbb{Z}_{q}^{N\times m}$ $\mathbf{S}\leftarrow_{R}\mathcal{D}$
$\mathbf{E}\longleftarrow_{R}\mathcal{X}_{\sigma}^{M\times m}$
$\mathbf{U}\leftarrow\mathbf{A}\mathbf{S}+\mathbf{E}$ $\mathbf{U}\leftarrow\mathbf{S}\mathbf{A}$
     $mpk\leftarrow(\mathbf{A},\mathbf{U}),\ msk\leftarrow\mathbf{S}$
[…]
procedure Enc($mpk,\mathbf{x}\in\mathbb{Z}^{m}$)
$\mathbf{r}\longleftarrow_{R}\{0,1\}^{M}$ $\mathbf{s}\leftarrow_{R}\mathbb{Z}_{q}^{N}$
$\mathbf{e}_{0}\leftarrow_{R}\mathcal{X}_{\alpha q}^{M}\ \mathbf{e}_{1}\leftarrow_{R}\mathcal{X}_{\alpha q}^{m}$
$\mathbf{ct}{{}^{\prime}}\leftarrow\mathbf{A}^{\top}\mathbf{r}$ $\mathbf{ct}{{}^{\prime}}\leftarrow\mathbf{A}\mathbf{s}+\mathbf{e}_{0}$
$\mathbf{ct}{{}^{\prime\prime}}\leftarrow\mathbf{U}^{\top}\mathbf{r}+t(\mathbf{x})$ $\mathbf{ct}{{}^{\prime\prime}}\leftarrow{\mathbf{U}\mathbf{s}+\mathbf{e}_{1}+\mathbf{x}\cdot\lfloor\frac{q}{K}\rfloor}$
     return $\mathbf{ct}\leftarrow(\mathbf{ct}{{}^{\prime}},\mathbf{ct}{{}^{\prime\prime}})$
procedure KeyGen($msk,\mathbf{y}\in\mathbb{Z}^{m}$)
$sk_{y}\leftarrow({msk\cdot\mathbf{y}},\mathbf{y})$ $sk_{y}\leftarrow({msk^{\top}\mathbf{y}},\mathbf{y})$
     return $sk_{\mathbf{y}}$
procedure Dec($\mathbf{ct},\ sk_{\mathbf{y}}=(\mathbf{d},\mathbf{y})$)
     $C\leftarrow\mathbf{y}\cdot\mathbf{ct}^{\prime\prime}-\mathbf{d}\cdot\mathbf{ct}^{\prime}\mod{q}$
     $res\longleftarrow\text{the plaintext }x\text{ that minimizes:}$:
${\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}|C-t(x)|}$ ${\color[rgb]{0.7890625,0.12109375,0.21484375}\definecolor[named]{pgfstrokecolor}{rgb}{0.7890625,0.12109375,0.21484375}|\left\lfloor\frac{q}{K}x\right\rfloor-C|}$
     return $res$
II-2 LWE-based FE
the LWE-based IPFE schemes that we selected are Bounded-Norm Inner Product schemes. This means that each plaintext $\mathbf{x}$ and vector $\mathbf{y}$ has to be respectively such that $\|\mathbf{x}\|_{\infty}<X$ and $\|\mathbf{y}\|_{\infty}<Y$ for some fixed $X,Y$.
The single-input selective-secure scheme is derived from Regev PKE and turned into a IPFE scheme as explained in [10], while the adaptive-secure single-input scheme is extensively presented in [8]. It is worth noting that the latter bases its security on a variant of the LWE problem called the multi-hint extended-LWE (mheLWE) problem. Although this variant can be seen as an instance of the LWE problem with a fixed number of samples and some additional information, it is not easier than the plain LWE problem, as there is a reduction from LWE to mheLWE [11].
Let $\mathcal{X}_{\sigma}$ denote an integer Gaussian distribution over $\mathbb{Z}_{q}$ with standard deviation $\sigma$ and $\mathcal{D}$ be a distribution over $\mathbb{Z}^{m\times M}$ as defined in [11]. Given two primes $p,q$ with $q>p$, for every $v\in\mathbb{Z}_{p}$ let the center function be defined as $t(v)=\lfloor v\cdot\frac{q}{p}\rfloor\in\mathbb{Z}_{q}$. Let $\alpha$ be a real number such that $\alpha\in(0,1)$. A brief description of both schemes is presented in Algorithm 2.
II-3 From Single-Input to Multi-Input
in order to lift the presented schemes to a multi-input setting, we used the compiler proposed by M. Abdalla et al. [8] and summarized in Algorithm 3. This compiler works when the single-input FE scheme satisfies two properties called Two-step decryption and Linear encryption.
Algorithm 3 $\mathcal{(MIFE)}$ Compiler of [8] that lifts a single-input IPFE $\mathcal{FE}$ to a multi-input setting.
procedure Setup($1^{\lambda},m,n$)
     for all $i\in[n]$ do
         $\mathbf{u}_{i}\longleftarrow_{R}\mathbb{Z}_{q}^{m}$
         $(mpk_{i}^{\prime},msk_{i}^{\prime})\longleftarrow\mathcal{FE}\text{.Setup}(1^{\lambda},m)$
         $csk_{i}\leftarrow(mpk_{i}^{\prime},\mathbf{u}_{i})$
     $msk\leftarrow((msk_{i}^{\prime})_{i},(\mathbf{u}_{i})_{i})$
     return $(msk,(csk_{i})_{i})$
procedure Enc($csk_{i},\ \mathbf{x}_{i}\in\mathbb{Z}_{q}^{m}$)
     $ct_{i}\longleftarrow\mathcal{FE}.\text{Enc}(mpk_{i}^{\prime},\mathbf{x}_{i}+\mathbf{u}_{i})$
     return $(ct_{i})_{i\in[n]}$
procedure KeyGen($msk,\ \mathbf{y}:\mathbf{y}=(\mathbf{y}_{i})_{i\in[n]},\ \mathbf{y}_{i}\in\mathbb{Z}_{q}^{m}$)
     for all $i\in[n]$ do
         $sk_{i,y}\longleftarrow\mathcal{FE}.\text{KeyDer}(msk_{i}^{\prime},\mathbf{y}_{i})$
     $z\leftarrow\sum_{i\in[n]}\langle\mathbf{u}_{i},\mathbf{y}_{i}\rangle\in\mathbb{Z}_{q}$
     $sk_{y}\leftarrow((sk_{i,y})_{i},z)$
     return $sk_{y}$
procedure Dec($sk_{y},\ ct_{1},\dots,ct_{n}$)
     for all $i\in[n]$ do
         $D_{i,1}\longleftarrow\text{Dec}_{1}(ct_{i},sk_{i,y})$
     $res\longleftarrow\text{Dec}_{2}\Bigl{(}\sum_{i\in[n]}D_{i,1},z\Bigr{)}$
     return $res$
When Federated Learning is combined with MIFE, an additional entity is often introduced: the Third Party Authority (TPA). This entity is responsible for setup, key generation, and key distribution. Additionally, the entity that performs the decryption is called aggregator.
III Related work
One of the main concerns in FL is the risk of data leakage due to various types of attacks. This privacy concern can be classified into two main categories[12]: privacy of the local model and privacy of the output model. Privacy of the local model means that no-one, including the aggregator, should have access to updates of the individual clients model. Privacy of the output model means that no one can extract information about the training data from the model computed by the Aggregator at each round.
A scheme that satisfies both privacy requirements is called Privacy-Preserving Federated Learning (PPFL) [13].
To obtain privacy of the local model some cryptographic primitives can be used. The main directions are represented by Homomorphic Encryption (HE), Multi Party Computation (MPC) with Secret Sharing techniques and FE. As pointed out in [13], there are use-cases in which FE approaches are the most promising. In this work we follow this direction by focusing on such primitive.
Regarding the privacy of the output model, a well known approach is Differential Privacy (DP) [12]: a non-cryptographic mechanism that consists in adding some noise to the information that are sent to the aggregator. A widely popular algorithm, that adds noise during the training process, is the Differential Private Stochastic Gradient Descent (DP-SGD).
HybridAlpha [7] represents the first example of the usage of MIFE in Federated Learning. The involved entities are the Aggregator, the Clients and the TPA that handles key generation and distribution. This protocol uses DP and MIFE to solve some privacy concerns, but leaving some security issues regarding the privacy of the local model. Building on the HybridAlpha framework, other works have explored developing PPFL using DP mechanisms and MIFE [14, 15, 16].
Other solutions have moved towards a decentralized setting in order to not rely on a TPA [13]. The main problem of these solutions is the need of cross-client interactions, which result in additional communication overhead and poor scalability.
IV Threat model
We consider a FL scenario with an honest-but-curious server that follows the correct procedures for parameter aggregation and client selection, but that may attempt to infer clients’ confidential information by exploiting the knowledge shared by clients during the federated training process. This information includes the model’s parameters and the global model’s accuracy on local validation sets (in the case of FLAD). Using this information, along with knowledge of the global model’s architecture, the server could attempt to reconstruct private information from the clients’ training data [5].
We assume that a subset of clients may be dishonest and collude to obtain private information from other clients (e.g. they could share their model to infer information on the other clients’ parameters). Throughout the training process, we suppose that at least two clients among the participants to the protocol are honest.
Moreover, we assume that the clients do not have the ability to manipulate the training data to compromise the global model’s operations. This type of attack, known as poisoning attack is a well-known problem [17, 18], but it is outside the scope of this work.
Finally, the TPA, which is responsible for distributing the keys for MIFE and the signatures needed to authenticate communications, is assumed to be honest by both the clients and the server.
Our work is based on HybridAlpha framework. HybridAlpha suffers from some security issues [13], in particular, there exists two attacks that leads the server to obtain more information about the clients than what it should.
The first attack, referred to as ciphertexts mix-and-match, decrypts sums of ciphertexts from different rounds to ob- tain other information. The second attack, called decryption- keys mix-and-match, allows the server to retrieve the model parameters sent by a client using different decryption keys from different rounds. To mitigate these leakages, the Multi-RoundSecAgg framework was introduced in [19]. This solution presents a trade-off between privacy and convergence time of the global model. Additional solutions have been proposed in a decentralized setting, as in [13]. Our scheme addresses these data leakages without using these approaches.
V Scheme description
In this section we describe our scheme in full detail, by listing all the interactions between the entities involved, in the framework of FLAD and using MIFE as a building block.
The scheme we propose, briefly described in Figure 1, is a modified version of FLAD [2], to which a layer of security is added by exploiting MIFE. Note that in this section we abstract from a specific MIFE instantiation (e.g. DDH, LWE, selective or adaptive).
First, the TPA performs the setup and generates the keys. Then, it distributes to each of the $n$ clients their secret key $csk_{i}$ and a common seed $s_{0}$. This seed, along with a previously agreed collision resistant pseudo-random function $f$, is used to generate some labels $\gamma_{t}=f^{t}(s_{0})$ that, added to the plaintexts are used to identify the rounds, as noted in [13]. Additionally, the TPA provides the server with the decryption key $sk_{y}$ associated with the vector $y=(1,\dots,1)\in\mathbb{Z}^{n}$. This is the only key the server needs, since in each round it has to compute the sum of the clients’ parameters and scores. Moreover, in our implementation we set the parameter $m$, that represents the number of entries of each plaintext vector that have to be summed, to $1$. This is because the server has to compute sums across corresponding entries of the input vectors of different clients. In fact, with $m>1$, the server would also compute the sum among the first $m$ entries of each client’s own input, consistently with 1, resulting in an output that is meaningless for our purposes.
At this point, the FLAD’s algorithm InitClients($\omega_{0},c_{s},c_{e}$) is run by the clients themselves – and not by the server as in the original scheme [2] – as every input parameter is public. This means that every client initializes its own model and for the first round everyone trains with the maximum amount of steps and epochs.
At each round, every client either performs local training or not, depending on the personal score obtained in the previous round. Regardless of whether training was performed, the client encrypts its parameters and sends them to the server. In the encryption phase the client computes the label of the current round, adds it to the plaintext and performs .$\mathcal{MIFE}.Enc$. Note that this step is done without batching.
The server computes .$\mathcal{MIFE}.Dec$ on the $n$ received ciphertexts, obtaining the sum of the parameters along with $n$ times the label that identifies the round. It then sends this result to each client, who subtracts the label and divides by $n$, to obtain the mean of the client’s parameters. Each client uses this mean to get the accuracy scores on their own validation set. Then, they encrypt their score using the same procedure followed for parameters’ encryption and send it to the server. The server computes .$\mathcal{MIFE}.Dec$ and returns the result to the clients, which – knowing the label – can compute the mean accuracy score value. The latter is used to evaluate the progress of the training and decide if the client will train in the next round. The algorithm that assigns steps and epochs for the next round to a client, in FLAD [2] makes use of the minimum and maximum of the clients’ accuracies for the current round. Since in this case the latter are unknown to the server, $\sigma$ is computed as $\lvert\frac{a^{\mu}-a^{i}}{a^{\mu}}\rvert$, as it is a natural alternative for the purpose. Note that steps and epochs for a specific client are computed by the client itself, as it holds both its $a^{i}$ and $a^{\mu}$. This is also good from a privacy perspective. We assume that during the training process, each client uses a differential privacy mechanism as in [7]. We will not dive further into it as it is out of scope for this work.
Figure 1: Protocol description.
V-A Termination Criterion
The FL process terminates following the same procedure of [2]. The difference is that the decision process is moved to the client side, as the server does not know the mean and maximum score. When the process concludes, the clients report it to the server. In case some clients are dishonest and tell to the server that the process is not ended, the TPA will intervene. Using the log saved by the server of the last $l$ mean scores decryption, where $l$ = patience, the TPA will be able to identify the dishonest clients.
V-B Joining and dropouts of clients
Our scheme also supports clients to join and dropout. Let $n$ be the number of clients currently participating to the process and $t$ the current round. These are public parameters known by all entities.
For security reasons, at each round, at least $2$ clients have to be honest. Consequently, if we allow up to $s$ drop out, we need to start with at least $s+2$ honest clients.
To avoid a fresh setup each time a new client joins the training, the TPA generates more keys in the first setup phase. When a client joins, it receives the private key and the seed by the TPA. The latter updates the decryption key $sk_{y}$ and sends it to the server. The amount of exceeding keys is strongly related to the specific use-case and is out of scope for this paper. However, it is worth noting that associating a new $u_{i}$ to the new client exposes our scheme to a mix and match attack (Section IV) on the updated $sk_{y}$. In fact, this new $u_{i}$ can be directly obtained subtracting $z$ in the old $sk_{y}$ to the new $z$. The simplest way to avoid this is changing the $u_{i}$ of a random client when a new client joins.
If a client $i_{0}$ drops out the process, for the scheme to keep on working with the remaining subset of clients, it suffices to update $sk_{y}$, by removing the term depending on $u_{i_{0}}$. Also in this case, changing a random $u_{i}$ among the remaining clients is needed to avoid a mix-and-match attack, similar to the user joining scenario.
We want to emphasize that this resulting scheme does not have the privacy leakages of HybridAlpha that [13] highlights. Indeed, an honest-but-curious server can not obtain more information about the clients than an honest aggregator would.
The ciphertexts mix-and-match attack is not applicable in our scheme due to the usage of the labels that identify a certain round. In fact, ciphertexts that come from different rounds can not be decrypted together.
Also the decryption-keys mix-and-match attack is infeasible due to the way in which FLAD works, since the decryption key is the same in all rounds. The situation in which new clients join or drop out the training process is more critical. Indeed, when the number of users varies, the private keys of individual clients remain unchanged while the decryption key changes. However, the use of labels and the change of one client’s $u_{i}$ protects our scheme from this attack.
VI Experimental setup
In our experiments, we used the latest version of FLAD’s code at the time of writing (Release 1.0 [20]), executed in a Python 3.9 environment with Tensorflow version 2.7.1. All experiments have been performed on a server-class computer equipped with an AMD EPYC 9454P 48-Core Processor and 128 GB of RAM.
[…]
Figure 2: Mean final accuracy and number of rounds relative to decimal digits truncation.
The experimental settings are consistent with those documented in the FLAD’s paper [2] in terms of dataset (CIC-DDoS2019 [21]), data preprocessing (an array-like representation of traffic flows, where the rows of the grids represent packets in chronological order and columns are packet-level features), unbalanced and heterogeneous data across the clients (one and only one attack assigned to each client) and ML model (a Multi-Layer Perceptron (MLP) with two hidden layers of $32$ neurons each, and an output layer with a single neuron for binary classification of the network traffic as either benign or DDoS). The total number of parameters of this model is 4 64146414\ 6414 641.
In order to achieve $128$ bits of security, we used Lattice Estimator [22] to learn the LWE parameters, along with the considerations in [11], and the NIST guidelines [23] for the DDH parameters.
Please note that all benchmarks and experiments are conducted using datasets available in the original FLAD repository [20]. We approximate floating points to integers by multiplying them for a common factor $10^{\Delta}$ and truncating the remaining digits. We set $\Delta=2$ since it is the most efficient approximation, achieving both small bitlength of the plaintexts and good performance in federated training, as in Figure 2. Data were collected on 11110 federated trainings with different random initial parameters, for each value of $\Delta$.
We also exploit parallelization to speed up both encryption and decryption of the models. We divided the plaintext (ciphertext, respectively) list into $15$ chunks with the same size. Then, we assigned a different process to each chunk.
$N$ $M$ $log(q)$ $\alpha$ $log(q)$
LWE 80 $5327$ $63$ $1.09\times 10^{-28}$ DDH 3072
38 $9462$ $248$ $1.71\times 10^{-53}$ 3072
TABLE I: Parameters to achive $\approx{128}$ bits of security.
VII Experimental results
In this section we discuss the memory cost and the efficiency of the proposed scheme, by running a full training. We use a different cryptographic protocol in each run, to compare their performance. As we are about to see, despite the fact that LWE-based approaches grant post-quantum security, their selective version copes with DDH in terms of time consumption, but it is highly memory-consuming.
In Table II we compare the proposed schemes according to the memory cost. This comparison does not take into account the labels impact. Note that, since the vector y consists entirely of ones, each entry requires only one bit for storage.
$\mathbf{|cks_{i}|}$ $\mathbf{|sk_{y}|}$ $\mathbf{|ct|}$
DDH $2\log{(q)}$ $(n+1)\log{(q)}+n$ $2\log{(q)}$
$3\log{(q)}$ $(2n+1)\log{(q)}+n$ $3\log{(q)}$
LWE $\log{(q)}(MN+M+1)$ $\log{(q)}(nN+1)+n$ $\log{(q)}(N+1)$
$\log{(q)}(MN+N+1)$ $\log{(q)}(nM+1)+n$ $\log{(q)}(M+1)$
TABLE II: Memory cost DDH-based and LWE-based schemes.
Plugging the parameters shown in Table I in the formulas in Table II, a notable difference is the space required for each client’s private key in DDH-based schemes respect to LWE-based ones. In particular, the formers require approximately $1$ KiB, whereas LWE-based schemes need $3$ to $10$ MiB. What stands out more about this comparison is the space required for each ciphertext in the adaptive LWE-based scheme, which is approximately $285$ KiB, whereas in the other three schemes it is around $1$ KiB. In our test, with $13$ clients and 4 64146414\ 6414 641 parameters, we get that at each round the ciphertexts use around $16.5$ GiB of memory. This shows the poor scalability of this particular scheme compared to the others.
Figure 3 shows a comparison between the average time needed by the slowest client to train their model in a round of FLAD, the average time that the client uses to encrypt the model parameters and the aggregation time needed by the server to decrypt the parameters of the output model. It should be noted that in both DDH-based schemes, the aggregation algorithm requires performing a discrete logarithm. We solved it by precomputing a lookup table.
Figure 3 shows that the encryption in DDH-based schemes is more efficient compared to LWE-based schemes. When considering the total time required for a full round, the selective LWE-based scheme outperforms the DDH-based scheme, while the adaptive LWE-based scheme is notably less efficient. It is important to note that LWE-based schemes offer quantum resistance, a feature not shared by DDH-based schemes.
DDHselectiveDDHadaptiveLWEselectiveLWEadaptive00202020204040404060606060808080801001001001001201201201203.213.213.213.213.193.193.193.193.23.23.23.23.33.33.33.311.3411.3411.3411.3414.7614.7614.7614.7622.0622.0622.0622.0632.2332.2332.2332.2348.3148.3148.3148.3193.9693.9693.9693.962.452.452.452.45108.26108.26108.26108.26Time (sec)FLAD Training TimeEncryption TimeAggregation Time
Figure 3: Time comparison of different phases within a round.
VIII Conclusion and Future Work
In this paper, we have addressed both cyphertexts mix-and-mach and decryption-keys mix-and-match attacks in the context of Federated Learning (FL) applied to network security problems. To achieve this, we have integrated MIFE techniques with a recent adaptive FL implementation for network intrusion detection, namely FLAD.
While preventing reconstruction attacks, our experimental result show that adding a layer of cryptography heavily impacts the memory and time performances of FLAD. This is clear when the primitives are istantiated with security parameters that are appropriate for the modern standards. However, these findings give a realistic benchmark on the performances of two standard choices such as DDH-based and LWE-based MIFE, and represent a major improvement compared to similar studies in the scientific literature.
Surprisingly, the selective version of LWE-based MIFE can actually outperform its DDH counterpart, while preserving a moderate memory consumption and being post-quantum secure. Regarding the adaptive LWE-based MIFE, our work shows that it is not usable in a realistic scenario. This is fostering a deeper study on different quantum resistant solutions, such as RLWE (see [24] for more information).
In a further study, we want to compare the performances of different post-quantum secure primitives and to test and propose new batching routines that allows to process the client’s plaintexts at once. Another interesting aspect we aim to investigate is optimizations of MIFE that further exploit the specific features of FLAD.
Acknowledgment
This work was supported by Ministero delle Imprese e del Made in Italy (IPCEI Cloud DM 27 giugno 2022 – IPCEI-CL-0000007) and European Union (Next Generation EU).